OpenStudy (anonymous):

HELP ASAP!!! WILL MEDAL!!!! Which step should be completed first to solve the equation -2 = 3z + 4? A. Add 2 to both sides. B. Subtract 4 from both sides. C. Divide both sides by 3. D. Add 4 to both sides.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

thanks can you help with more?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

can you help me with more? i'll follow you

OpenStudy (campbell_st):

Step B is the 1st step needed to solve the equation its a 2 step problem with the 2nd step be C

OpenStudy (anonymous):

oh ok can you help with more?

OpenStudy (anonymous):

soooo are you gonna help me with more or no?

OpenStudy (anonymous):

@campbell_st ?

OpenStudy (anonymous):

Yes

OpenStudy (britanyashton):

add two to each side. thats how i would start out with this problem.

OpenStudy (anonymous):

What is the solution to the equation 8 - 2b = -2/3(12b + 6)? b = ?

OpenStudy (britanyashton):

actually nvm thats if you want to graph. sorry!

OpenStudy (anonymous):

?

OpenStudy (anonymous):

B=-2

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

This web site might help you :) http://www.mathpapa.com/algebra-calculator.html

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

more?

OpenStudy (anonymous):

actually never mind lol thnx anyway!

OpenStudy (anonymous):

the two questions answered are correct!!!! YAY!!!